OpenStudy (konradzuse):

Carbon-14 dating assumes that the carbon dioxide on Earth today has the same radioactive content as it did centuries ago. If this is true, the amount of 14C absorbed by a tree that grew several centuries ago should be the same as the amount of 14C absorbed by a tree growing today. A piece of ancient charcoal contains only 45% as much of the radioactive carbon as a piece of modern charcoal. How long ago was the tree burned to make the ancient charcoal? (The half-life of 14C is 5715 years. Round to the nearest year.)

OpenStudy (konradzuse):

y = Ce^kt 1/2C = Ce^k(5715)

OpenStudy (konradzuse):

k = 1/5715 ln(1/2)

OpenStudy (konradzuse):

0.45C = Ce^[ln(1/2)/57145]t ?

OpenStudy (konradzuse):

ln(0.45) = ln(1/2)t/5715??

OpenStudy (chaise):

I'm not too sure on this one, sorry.

OpenStudy (konradzuse):

6584 I think...

OpenStudy (unklerhaukus):

that sounds like a good figure if the is a bit less then half the C-12 the sample should be a bit over 1 half life old